# Can LLM-Simulated Practice and Feedback Upskill Human Counselors? A Randomized Study with 90+ Novice Counselors

### v1

@@ Abstract
Figure 1: Our between-subjects study randomizes participants to either practice with AI Patients alone (P) or practice with AI patients and receive AI feedback (P+F). We holistically evaluate counselor skill development from three perspectives: automatic assessments of behaviors of skills used via LLM classifiers; self-efficacy and its calibration with actual performance; and qualitative self-reflections after the training intervention chat and post-intervention chat.
We conducted a 75-minute online lab study with novice counselors (N=94) to empirically evaluate how different LLM-simulated practice modes in CARE impact skill development. Participants were randomly assigned to one of two conditions: (1) Group P: Practice with LLM-simulated patients without AI feedback, or (2) Group P+F: Practice with LLM-simulated patients plus AI feedback (see Fig. 1). To comprehensively study learning outcomes, we measured changes from pre-intervention to post-intervention across multiple dimensions: actual behavioral performance (calculated via automatic analysis of counseling transcripts), participants’ self-efficacy (perceived confidence in counseling skills), and metacognitive awareness (captured through guided self-reflection prompts). Our study design specifically addressed: What changes occur after practice with an AI-simulated patient alone? How do these outcomes differ when participants also receive structured AI feedback?
Our findings reveal that practice with feedback significantly improves empathetic listening skills, while practice alone shows minimal improvement in these core competencies. Notably, both groups failed to improve the accuracy of participants’ self-efficacy assessments, highlighting the challenges involved in helping learners develop a well-calibrated understanding of their own performance. Through analysis of participants’ self-reflections, we found the P+F group internalized the importance of empathetic and active listening, but P participants continued to overly focus on solutions, albeit with increased information-gathering. These insights suggest that future LLM-based training systems must integrate feedback mechanisms to cultivate the empathetic stance fundamental to effective counseling.
Stages Skill Category Description
Table 1: Our analysis of skill development has a theoretical-foundation in Hill’s Helping Skills model [5]. We select a skill subset relevant for beginning counselors at the undergraduate and first-year graduate level [36]. These include microskills during the exploration and action stages; and macro skills that are applicable throughout the session. Hill’s insight stage, of which self-disclosure was the only relevant skill for basic counseling, was excluded from our primary analyses due to its infrequent occurrence in our data.
1 Results
This section presents findings from our investigation of how LLM-based practice affects counselors’ skills development. To understand the impact of simulated practice alone (P) and practice with feedback (P+F), we integrate evidence from three complementary data sources.
Behaviors of Skills Used: Training systems should help improve the use of skills that align with evidence-based counseling practice. Thus, we analyzed the skills used in transcripts and how they change between the pre-intervention and post-intervention assessments using high-accuracy LLM classification models. We assessed behaviors of skills used for the exploration stage (strong uses in empathy, reflections, questions) and action stage (suggestions needing improvement) of Hill’s Helping Skills framework [5]; see Table 1 for definitions.
Self-Efficacy Measures: It is informative to understand participants’ confidence and their awareness of their own strengths and limitations. We measured counselor self-efficacy using the Counselor Activity Self-Efficacy Scale [37], which covers confidence for exploration skills, action skills, and session management skills (Table 1). Our analysis tests whether the lowest performing individuals overestimate their abilities (aka the Dunning-Kruger effect [38]) and whether LLM-training interventions improve the accuracy of their self-assessments and raw self-efficacy scores.
Qualitative Data on Learner Takeaways: LLM-simulated training provides opportunities for experiential learning opportunities [39] whereby reflection on action [34] can support updates to one’s knowledge. We also asked participants’ open-ended self-reflections following the practice intervention and post-intervention assessment, and qualitatively coded their responses.
We integrate these three data sources to provide a holistic view of how AI-based practice—with and without AI feedback—influences changes in counselors’ use of skills, self-efficacy, and key lessons learned.
1.1 Changes in Counseling Skill Behaviors
Our primary analysis uses LLM-based behavioral assessments of transcripts to evaluate whether skill usage changes from pre-intervention to post-intervention for both the practice and practice-and-feedback groups. In a secondary analysis, we investigate how exposure to good examples of skills within the AI feedback may explain any changes we observe.
Change after P Change after PF Differences in
change after P vs. P+F
[…]
Empathy ($\uparrow$) -9.6 0.001 -0.52 5.4 0.117 0.23 15 0.001 0.72
Reflections ($\uparrow$) 1.4 0.391 0.18 3.7 0.034 0.32 2.3 0.323 0.2
Questions ($\uparrow$) 2.4 0.421 0.12 6.6 0.018 0.36 4.1 0.296 0.22
Figure 2: Changes in counseling behaviors following AI patient simulations alone (P) versus AI patient simulations with AI feedback (P+F). The plot displays bootstrapped means for pre-intervention and post-intervention interactions. The table presents statistical comparisons with corresponding effect sizes, with bolded values indicating significance after Benjamani-Hochberg correction [40] for the 21 planned comparisons (12 for behavioral changes and 9 for self-efficacy changes). Notably, P group experiences a significant drop in strong uses of Empathy (-9.6% change, $d=-0.52$), whereas P+F group’s use of Empathy is maintained and trends towards improvement. The P+F group also experiences noticeable improvements in Reflections (+3.7% change, $d=0.32$) and Questions (6.59% change, $d=0.36$)
We find that practice alone is not enough; feedback during practice is necessary to promote desirable counseling behaviors in empathetic and active listening. AI feedback during practice (P+F) led to improvements in Reflections (+3.6% change, $p=0.034$) and Questions (+6.59% change, $p=0.018$), and trended towards improvement in Validation (+3.6% change, $p=0.083$), Suggestions (-5.45% change, $p=0.057$) and Empathy (+5.37% change, $p=0.117$). In contrast, practice alone (P) led to only improvements in Suggestions (-5.85% change, $p=0.011$), but no improvements in Reflections, Questions, and Validation, and significantly worse expression of Empathy (-9.6% change, $p<0.001$). Pairwise comparisons between the two conditions were substantial and significant for Empathy (15% relative difference, Cohen’s $d=0.72,p<0.001$) and non-significant for others.
To better understand AI feedback’s role, we further analyzed whether the level of exposure to specific feedback during practice affected these changes. Our analysis quantifies the exposure to Good Alternatives during Practice (GAP), defined as the number of alternative responses in the AI-generated feedback that exhibited strong uses of a skill, normalized by chat length. We hypothesized that greater exposure to GAP for a skill would lead to stronger use of that skill. To test this, we fit a lagged-linear regressor that predicts post-chat behaviors, controlling for pre-chat behaviors, with exposure to good alternatives ($T_{GAP}$) as a scalar predictor ($b_{1}=\beta_{0}+\beta_{1}b_{0}+\beta_{2}T_{GAP}$).
[…]
Figure 3: Counselor Self Efficacy (perceived ability to use skills) for participants grouped by behaviors of skills used (actual performance). Notes: Gaps depict miscalibration between actual and self-assessed percentile of performance for quartile groups with bootstrapped 95% CIs. We only visualize for data collected in the post-intervention.
1.2 Self-Efficacy and its Miscalibration with Behaviors
Our investigation of self-efficacy involved three types of analyses. First, we examine the calibration of self-efficacy ratings with actual performance and evaluate whether P or P+F interventions improve this. Second, we tested for significant changes in raw self-efficacy scores after practice or practice-and-feedback.
Our analysis comparing self-efficacy ratings with actual performance across skill quartiles finds support for the Dunning–Kruger effects more substantially for action skills and to a lesser degree for exploration skills. The interaction between measure and quartile was significant in four out of six ANOVAs for action skills, while only one out of six was significant for exploration skills (Table 4). Pairwise comparisons also showed a pattern indicative of a Dunning–Kruger effect (see Table 5, Table 6 and Fig. 3): People in the lowest quartile overestimated themselves the most. Those in the highest quartile—and to a lesser degree also those in the second-to-highest quartile—tended to underestimate themselves.
Change after P Change after PF Differences in
change after P vs. PF
[…]
Exploration Skills 0.36 0.004 0.44 0.13 0.338 0.14 -0.21 0.238 -0.25
Action Skills 0.27 0.166 0.21 0.33 0.026 0.34 0.04 0.884 0.03
Session Management 0.35 0.021 0.35 0.36 0.011 0.39 0.01 0.955 0.01
Figure 4: Changes in raw-scores of self-efficacy following AI patient simulations alone (P) versus AI patient simulations with AI feedback (PF). The plot displays bootstrapped means for pre-intervention and post-intervention. The table presents statistical comparisons with corresponding effect sizes, with bolded values indicating significance after Benjamani-Hochberg correction [40] for the 21 planned comparisons (12 for behavioral changes and 9 for self-efficacy changes).
Participants’ Ability to Self-Assess Their Skill Level Remained Mixed After LLM Practice. For the practice only (P) group, the mean discrepancy in exploration skills changes from 11.6 percentile underconfidence to 5.7 percentile overconfidence, a significant shift ($p<0.001,d=0.58$). Besides this, we found no other significant changes in calibration. No significant differences was found among the P group for discrepancy in action skills ($p=0.227,d=0.18$). The P+F group showed no significant calibration changes for exploration skills ($p=0.479,d=-0.10$) or for action skills ($p=0.393,d=0.13$). Finally, between-group differences were not significant for discrepancy in exploration skills ($p=0.191,d=0.27$) or action skills ($p=0.743,d=0.07$).
We do find modest overall increases in self-efficacy after P and P+F interventions, with different patterns of improvement across skills (Fig. 4). For the P group, confidence in exploration skills showed a significant increase ($0.36$ points on an 8 point scale, $d=0.44,p=0.004$). Confidence in session management skills showed a significant increase for the P+F group ($0.36$ points, $d=0.39,p=0.011$). While session management skills for the P group also trended towards improvement, it was not significant after correcting for multiple hypothesis testing ($0.35$ points, $d=0.35,p=0.021$). Similarly, while confidence in action skills for the P+F group also increased, this result was not significant after correction of multiple hypothesis tests ($0.33$ points, $d=0.34,p=0.026$). Finally, we found no significant differences between participants who received AI feedback (P+F) versus those who did not (P) (across the three self-efficacy subscales, $d=-0.25,0.03,0.01$, $p=0.238,0.884,0.955$).
1.3 Qualitative Self-Assessments after AI practice with and without AI feedback
To understand participants’ learning process during CARE practice, we analyzed qualitative self-reflections from two time points: immediately after the training intervention chat, where participants responded to “What would you do differently as a therapist?” and after the post-intervention chat, where they reflected on “What did you do well as a therapist?”. We examined how initial intentions translated into reported strengths across the P+F and P groups. Two high-level themes emerged in how participants approached skill improvement:
(1) The P+F group expressed greater intentions and successes in improving their use of empathy and listening skills. P+F participants reported effectively using empathy (27%), validation (27%), and open-ended questions (52%). They emphasized the value of listening skills, such as reflective responses to signal understanding: “I should rephrase what they say so they know I’m understanding them” (P51). They also recognized that counseling should support client exploration of thoughts and emotions, rather than provide direct solutions. One participant reflected on this shift: “I asked them to expand on their feelings, rather than guiding them to my idea” (P39). AI feedback encouraged this shift toward providing emotional support and fostering client autonomy, helping counselors adopt a more empathetic and client-centered approach.
(2) The P participants remained solution-oriented but changed their approach to first gather information. While P+F participants intended to give fewer suggestions, many P participants continued to view suggestions as a central skill. In the post-intervention, 48% of P participants reported using suggestions successfully, compared to only 14% of P+F participants. Many P participants justified their continued use of suggestions by citing a desire to provide tangible, actionable help, for example, “Maybe because I’m untrained and solution oriented. I do not want to leave them with nothing, and nowhere to go” (P40). Some reflected on modifying how they delivered suggestions, emphasizing strategies like gathering more information to tailor advice or providing more concrete guidance. However, in the absence of feedback, most remained fixed in their approach, with some even reporting efforts to rephrase the same solution repeatedly to persuade the client.
2 Discussion
Our findings demonstrate that teaching counselors to implement client-centered approaches requires more than just simulated practice opportunities—it requires targeted feedback that guides novices away from their natural solution-oriented tendencies. The fully-featured version of CARE–combining LLM-simulated patient practice and LLM-based feedback–helped participants improve their use of empathetic and active listening skills, with notable improvements in questions ($d=0.36$) and reflections ($d=0.32$). In comparison, practice with an AI patient alone only led to fewer inappropriate suggestions ($d=-0.39$), but no improvements in reflections or questions, and significantly worse uses of empathy (-9.6% change, $d=-0.52$; 15% relative difference to P+F, $d=0.72$). These effect sizes are comparable to those found in studies of human supervision during standardized roleplays, where Maaß et al. (2025) [41] reported observer-rated skill improvements with effect sizes ranging from $d=0.29-0.49$. Since LLM-simulated practice and feedback are not bottlenecked by the resource constraints of human trainers, AI training systems like CARE show promise in scaling access to effective counseling training.
While the combination of feedback and simulated practice was key to CARE’s success, we also recognize that CARE represents just one approach to providing expert-aligned LLM feedback to learners. CARE allows trainees to request feedback on responses they have already sent. This approach allows trainees to experience productive learning discomfort by trying and making mistakes in simulation, while still offering feedback upon request to refine future responses. This design choice parallels elements of both live human supervision [42] and delayed supervision after a session [41], creating a hybrid approach that may offer flexibility based on trainee preferences. Given the promise of feedback combined with practice, the AI for psychotherapy field should evaluate different AI-feedback designs against each other. Several other possible strategies for giving feedback include just-in-time skill suggestions [21, 43], counterfactual simulations [44], and global session feedback. While conducting randomized trials in medical education research can be tricky [45], randomized trials delivering different training interventions to large participant-pools for longer durations [46] could provide essential data to determine optimal approaches for different learning objectives and learner characteristics.
Our study revealed that participants’ self-efficacy ratings were poorly calibrated with their actual performance, especially among lower performers. This finding, consistent with prior research, suggests that self-assessment accuracy alone may not be a reliable indicator of counselor competence or development. Recent reviews indicate that efforts to improve self-assessment calibration have limited impact on learning or performance outcomes [47]. Instead, it is valuable to objectively assess specific standards of performance and skill use [42, 48] and design interventions that can help low performers improve on those metrics while maintaining a positive morale for continued practice. As AI-based training tools evolve, integrating objective performance measures and structured self-reflection, rather than relying solely on self-assessment, offers a more robust approach to supporting counselor development.
Several limitations of this study warrant acknowledgment. First, our assessment focused primarily on immediate skill acquisition rather than long-term retention or transfer to actual clinical settings. Second, while our automated behavioral analysis focused on microskill usage, it could not capture observable session-level characteristics. While traditional studies have employed human observers to provide such ratings, recent AI research has shown the validity of using finetuned LLMs in certain contexts to approximate session-level measures, such as working alliance [49], which might enable scalable behavioral analyses of more skill development constructs. Third, we recruited from a sample of bachelor-level counselors through the online Prolific platform, which may differ from in-person or classroom settings. For example, existing training environments for counselors will typically offer didactic materials (e.g., lectures and textbooks) and non-AI training opportunities (e.g., peer-to-peer roleplay and feedback). Thus, future work should evaluate the utility and acceptability of AI-based training tools within existing educational and training pathways. Finally, the efficacy of LLM-based practice and feedback training was only shown for client-centered microskills. Future work is required to understand whether these results generalize to other therapy modalities (e.g., psychodynamic, CBT/ACT), where training interventions may work to shift natural tendencies in similar or subtly different ways.
3 Conclusion
Despite increasing interest in using LLMs in mental health, to our knowledge, this is the first study to conduct a large-scale evaluation ($N=94$) of an LLM-based training system for developing core skills in novice counselors. Our findings show that while simulated practice (P) proved insufficient—and in the case of empathy, potentially harmful—practice with AI-generated feedback (P+F) supported measurable improvements in key counseling skills of demonstrating empathy, delivering reflective responses, and engaging in client-centered inquiry. By combining realistic patient simulations with expert-aligned, skill-specific feedback, the CARE AI Training system can help novices to master skills involved in client-centered therapy, offering a promising avenue for scalable, evidence-aligned training in mental health care.
4 Methods
4.1 Overview of CARE Training Platform

### v2

@@ 1. Introduction
Figure 1. Our experiment randomizes participants to either practice with AI Patients alone (P) or practice with AI patients and receive AI feedback (P+F). We holistically evaluate counselor skill development from three perspectives: automatic assessments of behaviors of skills used via LLM classifiers; self-efficacy and its calibration with actual performance; and qualitative self-reflections after the training intervention chat and post-intervention chat.
Figure illustrating the randomized pre-post study design of CARE, an LLM-based counselor training system. Ninety-four novice counselors first completed a 5-minute tutorial and a 10-minute pre-intervention chat with an AI-simulated patient. Participants were randomized into two groups: Practice-only (P), who engaged in a 20-minute chat with an AI patient without feedback, and Practice+Feedback (P+F), who engaged in the same practice while receiving structured AI feedback, labeling strengths, weaknesses, and alternative responses across core microskills (empathy, reflections, questions, suggestions). All participants then completed a fixed 10-minute post-intervention chat with a new AI patient. Skill development was assessed across three dimensions: (1) behavioral performance, via fine-tuned LLM classifiers detecting strong or weak uses of counseling skills; (2) counseling self-efficacy, using the CASES-R survey of exploration, action, and session-management skills; and (3) qualitative self-reflections, where participants reported what they would do differently and what they did well. Figure illustrating the randomized pre-post study design of CARE, an LLM-based counselor training system. Ninety-four novice counselors first completed a 5-minute tutorial and a 10-minute pre-intervention chat with an AI-simulated patient. Participants were randomized into two groups: Practice-only (P), who engaged in a 20-minute chat with an AI patient without feedback, and Practice+Feedback (P+F), who engaged in the same practice while receiving structured AI feedback, labeling strengths, weaknesses, and alternative responses across core microskills (empathy, reflections, questions, suggestions). All participants then completed a fixed 10-minute post-intervention chat with a new AI patient. Skill development was assessed across three dimensions: (1) behavioral performance, via fine-tuned LLM classifiers detecting strong or weak uses of counseling skills; (2) counseling self-efficacy, using the CASES-R survey of exploration, action, and session-management skills; and (3) qualitative self-reflections, where participants reported what they would do differently and what they did well.
To address this gap, we develop CARE by combining two previous methods for realistic patient creation (Louie et al., 2024) and feedback generation (Chaszczewicz et al., 2024), into a wholistic system for training counseling skills, and conduct a randomized experiment to investigate how two different modes of simulated training impact counselor skill development. CARE enables (1) realistic practice with LLM-simulated patients, whose prompts are seeded by expert counselors to resemble challenging behaviors (Louie et al., 2024), and (2) structured feedback from a fine-tuned LLM that identifies strengths and areas for improvement across core counseling skills (e.g., empathy, reflections, questions, suggestions), while also providing explanatory rationale and alternative responses (Chaszczewicz et al., 2024). The generated feedback is grounded in established counseling frameworks (Nelson-Jones, 2013; Miller et al., 2003) and informed by expert-annotated examples, ensuring alignment with recognized training standards. Uniquely, CARE’s LLM patients and feedback have been co-designed, iteratively improved, and rigorously validated by counseling domain-experts, ensuring that our experiment on novice skill development controls for the quality of the LLM components’ outputs.
We conducted a 75-minute online lab study with 94 novice counselors to evaluate how different LLM-simulated practice modes in CARE impact skill development. Participants were randomly assigned to one of two conditions: (1) Group P: Practice with LLM-simulated patients without AI feedback, or (2) Group P+F: Practice with LLM-simulated patients plus AI feedback (see Fig. 1). We measured changes in behavioral performance (via transcript analysis), self-efficacy (via survey items), and intentions for growth (via self-reflection prompts). Our study design specifically addressed: What changes occur after practice with an AI-simulated patient alone? How do these outcomes differ when participants also receive structured AI feedback? Our results show the practice-and-feedback group significantly improved in their use of reflections and questions (d=0.32-0.39, p<<0.05), and trended toward improvement in empathy (d=0.23) and suggestions (d=-0.28). In contrast, the practice-only group only showed significant improvements in suggestions (d=-0.39), but actually worsened in Empathy (d=-0.52, p=0.001). Between-group comparisons show a substantial advantage for the P+F group in Empathy (d=0.72, p=0.001), indicating a strong feedback effect. In contrast, suggestions showed near-zero between-group differences (d=0.02, p=0.910) despite pre-post improvements in both conditions, suggesting that another mechanism besides feedback is driving this change. Through qualitative analysis of participants’ self-reflections, we found that the practice-and-feedback group internalized the importance of empathetic and active listening; however, practice-only participants continued to overly focus on solutions, albeit with increased information-gathering. Our discussion details possible reasons for skill changes in the practice-alone group: novice counselors evolve therapeutic intentions from observable patient behaviors–when the AI patients consistently expressed skepticism to suggestions, counselors reduce inappropriate use of suggestions; but when AI patients show no differential response to empathetic vs. non-empathetic statements, the P group gradually de-prioritizes empathy. Together, these results suggest that LLM-simulated training should integrate structured feedback to cultivate a client-centered, empathetic listening approach fundamental to effective counseling.
In summary, we contribute: (1) the design of CARE, an LLM-based training system that integrates realistic patient simulations with structured feedback grounded in counseling frameworks; (2) evidence from a randomized evaluation with 94 novice counselors, triangulating outcomes across behavioral performance, self-efficacy, and therapeutic intentions; and (3) design implications for LLM-simulated training, showing how structured feedback prevents empathy decline and supports effective counselor development, while highlighting ongoing challenges in improving overall self-efficacy while minimizing mis-calibration with performance.
2. Related Work
Our work training novice counselors via LLM-based systems is grounded in two areas of work. First, prior training approaches for clinical and communication skills have long relied on simulated patients, ranging from human role-plays to scripted virtual patients and, more recently, LLM-based simulations with automated feedback. Second, prior HCI research evaluating human-AI systems, especially in domains like health and education, emphasizes not only AI system accuracy but also how users learn, calibrate, and reflect when interacting with AI systems.
2.1. Training Systems for Clinical and Communication Skills
Traditional training for helping skills—empathy, active listening, and communication—uses theory, expert demonstration, role-play, supervised practice, and experiential learning (Hill, 2020; Hill and Kellems, 2002; Martin and Johnson, 2024; Huerta-Wong and Schoech, 2010). These approaches are effective but hard to scale: peers and supervisors require coordination, and trainees can pick up unhelpful habits without oversight (Atkins et al., 2014). Simulated standardized patients (trained actors) are common in health education; meta-analyses show they improve communication, knowledge transfer, and confidence (Shin et al., 2015). Counseling training likewise uses peer role-plays and standardized scenarios to practice skills like reflective listening and empathy (Kühne et al., 2020; Ay-Bryson et al., 2023), yet such exercises remain resource-intensive and limited in availability.
Virtual patients
Virtual patient (VP) simulations—computer or embodied agents—recreate clinical encounters to train history taking, nonverbal communication, empathy, and counseling (Albright et al., 2016; Schoenthaler et al., 2017; Rein et al., 2018). They provide safe, repeatable practice without actors and have been applied to suicide prevention, adolescent substance-use screening, and antibiotic conversations (Rein et al., 2018; Burmester et al., 2019; Schoenthaler et al., 2017). For example, Murali et al. (2022) used a conversational agent to teach vaccination-related counseling to laypeople; commercial platforms like Skillsetter follow deliberate-practice models (Skillsetter, 2024). Early VP systems were often scripted and template-based, costly to develop, and typically limited to single cases, reducing realism and diversity (T., 2003; Han et al., 2015; Othlinghaus-Wulhorst and Hoppe, 2020).
Early AI-based training systems
Building on VPs, early AI-driven systems experimented with automatically analyzing communication features and providing learners with feedback. For example, EQClinic visualized audio and video signals to help trainees reflect on their nonverbal behaviors in telehealth role-plays (Liu et al., 2016), while ConverSense detected and displayed social signals such as dominance and warmth from patient-provider interactions (Bedmutha et al., 2024). These systems raised self-awareness of communication styles, but their feedback was often decontextualized and difficult to apply in practice. Moreover, they did not directly target the counseling microskills important for effective therapeutic interactions.
LLM-simulated patients for role-play practice
Thus, a growing body of work in NLP and HCI has used LLMs to create simulated patients as role-play partners for counselor training (Wang et al., 2024; Louie et al., 2024; Stapleton et al., 2023; Steenstra et al., 2025). The goal of these systems is to provide practice environments that resemble real clinical encounters, making training more transferable and faithful to practice (Alinier and Oriot, 2022). However, achieving authentic simulations remains challenging. LLMs are highly sensitive to prompting (Zamfirescu-Pereira et al., 2023), and naive prompts in mental health contexts often produce unrealistic behaviors, including caricature, bias, and limited domain fidelity (Cheng et al., 2023).  Chen et al. (2023) found that naively prompting GPT-3.5 to simulate a patient profile with depressive symptoms led the chatbot to describe its emotions in formal, diagnostic language, which expert clinicians noted as inauthentic. Recent work prompts LLMs with psychology-grounded frameworks to simulate patients (e.g., Patient-Psi for CBT case conceptualizations (Wang et al., 2024) and multi-stage pipelines modeling cognitive factors (Steenstra et al., 2025)). Nonetheless, LLMs remain prompt-sensitive and typically need expert validation to capture realistic resistance, ambivalence, and other clinically relevant behaviors. We therefore adopt the expert-driven, iterative behavioral refinement method of Louie et al. (2024)—integrating those validated principles into CARE to produce more authentic simulated patients.
Automatic scoring and feedback for counselor transcripts
A parallel line of work has developed automated methods to help peer counselors improve their skills. Scoring-based systems (e.g., ratio of questions to reflections in a transcript) provide metrics, but these approaches offer limited, actionable guidance on how to improve. By contrast, suggestion-based systems generate or rewrite candidate responses to model more effective behaviors. Research in clinical NLP has produced numerous models for classifying and scoring counseling transcripts (Tanana et al., 2016; Pérez-Rosas et al., 2019; Schwalbe et al., 2014; Huang et al., 2018; Yang et al., 2024; Min et al., 2022). Many focus on a single microskill, such as reflections, providing numeric feedback on usage frequency (Can et al., 2012; Pérez-Rosas et al., 2017; Shen et al., 2020; Min et al., 2022). Others examine skill distributions more broadly and their relationship to conversational success (Wang et al., 2023). While valuable for large-scale analysis, these approaches rarely translate into actionable feedback for scaffolding trainee learning. To make feedback more interactive, researchers have explored real-time rewriting and suggestion systems. For example, Saha et al. (2022) and Sharma et al. (2021) proposed response rewriting methods to enhance empathy. With the goal of increasing interactiveness,  Sharma et al. (2023) proposed HAILEY, a tool that modifies peer supporters’ responses, while  Hsu et al. (2025) generated strategy-aligned suggestions during live conversations. Although promising, studies show that just-in-time suggestions can distract learners and foster overreliance, sometimes leading to negative learning effects when AI support is withdrawn (Abbas et al., 2024; Bastani et al., 2025; Kazemitabaar et al., 2024).
While this previous work developed NLP models for specific counseling tasks, the ability to use LLMs as zero-shot or few-shot reasoners has enabled further research in this area. Nonetheless, naively prompting LLMs in a mental health context can lead to generated outputs that are characteristic of low-quality therapy (Chiu et al., 2024). Therefore, a training system that uses LLMs to generate feedback for counselors needs to take measures to ensure the outputs are faithful and robust, lest it teach or promote bad practices (Moran et al., 2025). Chaszczewicz et al. (2024) co-designed a feedback dataset with therapy supervisors and fine-tuned an open-weight LLM to produce explanatory, actionable feedback. CARE adopts this expert-validated, fine-tuned model. While other training systems have employed prompting-based approaches with coding manuals and few-shot examples (Steenstra et al., 2025), our approach differs by integrating expert feedback directly into model weights through fine-tuning on expert-annotated examples, rather than relying on in-context learning alone.
2.2. Evaluating Human-AI Systems
Evaluating human-AI systems requires more than assessing model accuracy or output quality  (Blodgett et al., 2024). In HCI and education research, effectiveness is judged by its impact on learners: how people acquire skills, calibrate their understanding, and integrate feedback into practice. This framing is important in counseling training, where evaluation concerns not only usability but also the development of interpersonal behaviors in sensitive, high-stakes domains.
Recent work highlights the limitations of traditional benchmarks, which often fail to capture generative model capabilities (McIntosh et al., 2026). This calls for dynamic and human-centered evaluations (Liao and Xiao, 2023; Elangovan et al., 2024; Khullar et al., 2025), that move beyond static model metrics and consider human outcomes, and are relevant when assessing interactive training systems. Thus, when we evaluate human-AI systems, additional challenges arise. Researchers must account for both the technical performance and also user impact (Weidinger et al., 2023). While guidelines exist for designing human-AI systems (amershi2019guidelines; Wright et al., 2020), less work addresses how they should be evaluated. Some frameworks capture process and user preferences in human-LLM interaction (Lee et al., ), others focus on safety (Weidinger et al., 2023) or domain-specific contexts (Lee et al., 2024). Tools such as SPHERE propose multi-dimensional evaluation cards to structure study design and improve transparency, but consensus on evaluation practices remains limited (Zhao et al., 2025).
In counseling contexts, these gaps surface in three ways. First, evaluation must triangulate across behavioral outcomes, self-efficacy, and learning, aligning with evidence-based psychotherapy work in deliberate practice (Ericsson and others, 2006; Schön, 2017). Second, calibration is critical: learners often misjudge their own performance (Kruger and Dunning, 1999; Eva and Regehr, 2005), and AI feedback may inflate confidence without improving skills (Macnamara et al., 2024). Third, user perceptions of realism, trust, and workload shape adoption: relational agent studies show that authenticity fosters engagement (Mallik and Gangopadhyay, 2023), while trust research highlights risks of distraction and overreliance (Glikson and Woolley, 2020; Buçinca et al., 2021). Finally, in sensitive domains, evaluation must weigh ethical and pedagogical guardrails: ensuring feedback preserves learner agency and avoids harmful or misleading guidance.
Taken together, evaluating human-AI systems requires a multi-dimensional perspective that integrates skill outcomes, calibration, perceptions, and responsible design. Yet, few studies have examined how LLM-based training systems affect novice counselors across these dimensions. Our work contributes by combining objective performance measures, self-efficacy surveys, and qualitative reflections to provide a holistic evaluation of LLM-driven counseling training.
3. CARE Training System